OpenStudy (anonymous):

Find the solution of the linear equation. 2/3 (6 - 3x) = (1 - x) + (3 - x) A. x = -8 B. x = 8 C. infinitely many solutions D. no solution Can someone explain this to me without giving me the answer? I need it step by step. :p

myininaya (myininaya):

First do you know what to when you have a(b+c)? Do you know the distributive property?

OpenStudy (anonymous):

Yes.

myininaya (myininaya):

So do that for the left hand side of the equation.

myininaya (myininaya):

let me know what you get after performing the property

OpenStudy (anonymous):

2/3 6 + 2/3 3x?

myininaya (myininaya):

\[a(b+c)=ab+ac\] \[\frac{2}{3}(6-3x)=\frac{2}{3}(6)-\frac{2}{3}(3x)\] I think there was a minus sign in between.

myininaya (myininaya):

you can write a(b-c) as ab-ac

myininaya (myininaya):

Now you need to simplify: \[\frac{2(6)}{3}-\frac{2(3)x}{3}\]

myininaya (myininaya):

what is 6 divided by 3?

OpenStudy (anonymous):

2

myininaya (myininaya):

\[2(2)-\frac{2(3)x}{3}\] so instead of 6/3 I will put 2 now what is 3/3?

OpenStudy (anonymous):

1

myininaya (myininaya):

\[2(2)-2(1)x\]

myininaya (myininaya):

ok instead of 3/3 I put 1 now can you simplify that from there?

myininaya (myininaya):

mainly you just need to do 2(2) and 2(1) now

OpenStudy (anonymous):

How do I do a(b)?

myininaya (myininaya):

If you can do division you should be able to do the multiplication part. But let me see if I can explain multiplication... It is just basically repeated addition. 2(2) means you have 2 groups of 2 or 2+2 which means you have 4 altogether another example: Pretend we have 5(8) this means we have 5 groups of 8 (or you can even see it as 8 groups of 5) Now let's go with the 5 groups of 8 that means we have 8+8+8+8+8 which is 40. Do you need more help with multiplication? Or more examples?

myininaya (myininaya):

I found it helpful and made me multiply bigger numbers faster by remembering my time tables.

myininaya (myininaya):

So 2(2)-2(1)x=?

myininaya (myininaya):

remember 2(2) means 2+2 and 2(1) means 1+1

OpenStudy (anonymous):

You mean like 2^2 and 1^2?

myininaya (myininaya):

2^2 means 2(2) 1^2 means 1(1)

myininaya (myininaya):

2^2 means 2(2) which also means 2+2 1^2 means 1(1) which also means 1

myininaya (myininaya):

but we had 2(1) which means 1+1 which is 2

myininaya (myininaya):

If you are doing this type of problem, you should have had some experience with multiplication.

myininaya (myininaya):

Are you sure you don't know how to multiply?

myininaya (myininaya):

http://www.study-skills-for-all-ages.com/image-files/multiplication_table_1.gif Here is a multiplication table. I'm not sure if you ever seen one of these. This makes multiplication faster remembering this table.

OpenStudy (anonymous):

I get that now. :) Can you just show me which part to do first? Can you do it step by step with the drawing tool so I can SEE how to do instead of reading it? I think that's what confuses me. Here's the problem again so I don't have to keep scrolling up. :) Find the solution of the linear equation. 2/3 (6 - 3x) = (1 - x) + (3 - x) A. x = -8 B. x = 8 C. infinitely many solutions D. no solution

myininaya (myininaya):

we were doing \[\frac{2}{3}(6-3x)\] we distributed \[\frac{2}{3}(6)-\frac{2}{3}(3x)\] then you said 6/3=2 and 3/3=1 so we got 2(2)-2(1)x then I asked you to simplify this part which is what again?

myininaya (myininaya):

Come on I already told you what 2(2) and 2(1) is you should be able to simplify 2(2)-2(1)x

myininaya (myininaya):

I'm not going to just do the problem for you. I don't see how that is helpful to you for me to just give you the answer.

myininaya (myininaya):

I said earlier 2(2)=4 and 2(1)=2 so 2(2)-2(1)x can be written as 4-2x so the left hand side of the equation can be written as 4-2x so the initial equation can be written as 4-2x=(1-x)+(3-x) I would like to see your attempt on at least simplify the right hand side

myininaya (myininaya):

the right hand side being (1-x)+(3-x)

OpenStudy (anonymous):

2(2)-2(1)=1 Because 2(2) = 2+2 2(1) = 1+1 I know

OpenStudy (anonymous):

I'm probably annoying you with this, I'm sorry. :) Now how do I add the x in, where you said 2(2)-2(1)x?

myininaya (myininaya):

2(2)=4 2(1)=2 so 2(2)-2(1)x is 4-2x I already simplified the left hand side I was asking you to do the right hand side

myininaya (myininaya):

Even if I asked you to do 2(2)-2(1) this is not equal to 1 But I didn't ask you to do 2(2)-2(1) I asked you to do 2(2)-2(1)x

myininaya (myininaya):

2(2)-2(1) is 4-2 and since 4 and 2 are like terms 4-2 can be simplified more 4-2 is 2 2(2)-2(1)x is 4-2x nothing else can be done 4 and 2x are not like terms